# Agile Temporal Discretization for Symbolic Optimal Control

### v1

@@ Abstract
As control systems grow in complexity, abstraction-based methods have become essential for designing controllers with formal guarantees. However, a key limitation of these methods is their reliance on discrete-time models, typically obtained by discretizing continuous-time systems with a fixed time step. This discretization leads to two major problems: when the time step is small, the abstraction includes numerous stuttering and spurious trajectories, making controller synthesis suboptimal or even infeasible; conversely, a large time step may also render control design infeasible due to a lack of flexibility. In this work, drawing inspiration from Reinforcement Learning concepts, we introduce temporal abstractions, which allow for a flexible time-step. We provide a method for constructing such abstractions and formally establish their correctness in controller design. Furthermore we show how to apply these to optimal control under reachability specifications. Finally we showcase our methods on two numerical examples, highlighting that our approach leads to better controllers.
@@ I INTRODUCTION
Due to the increasing complexity of modern engineered dynamical systems, controllers become more and more arduous to design. In the litterature, this paradigm shift has been coined by the academic and industrial communities as the cyber-physical revolution [13, 3, 15]. On the other hand, it is crucial to provide safety and/or performance guarantees for such systems [14]. In particular, we focus on reachability analysis [2], which consists in, starting from a certain subset of the state space, designing a controller to reach a given set while avoiding obstacles. Being able to solve such problems for complex dynamical systems is crucial in many applications (see e.g. [1] and [9] for applications to respectively autonomous driving and robotic aerial vehicles).
In this context, abstraction-based methods are promising tools to provide formal reachability guarantees [20], whereby a finite-state mathematical object (the abstract system) approximates the original dynamical system (the concrete system). Every abstraction-based control method follows the same systematic procedure:
@@
    Abstraction – The concrete system is approximated by a finite-state mathematical object;
@@
    Concretization – From the discrete controller, a concrete controller is refined for the original dynamical system.
@@
Continuous-time deterministic dynamical systems are typically abstracted with a finite state automaton as follows. The continuous state space is partitioned in a finite set of cells. The input space is discretized in a finite number of inputs, and a discrete-time version of the original dynamics is considered, where every step corresponds to a time interval $\tau$ in the concrete system. The states of the automaton correspond to the cells of the partitioned state space, and a transition between $s_{1}$ and $s_{2}$ with input $u$ is added if the set of reachable states from the cell $s_{1}$ intersects with the cell $s_{2}$. In this work, we focus on the time discretization. Indeed, in the context of reachability analysis, fixing a too small constant timestep $\tau$ leads to conservatism for at least two reasons:
@@
    Spurious trajectories – At each step of a trajectory in the abstract system, the information of where is exactly the state within the cell is lost. As a consequence, an error is propagated with the number of steps, and more and more trajectories are valid in the abstraction but cannot happen in reality: such trajectories are called spurious trajectories. This phenomenon is also illustrated in red (two small timesteps) and in blue (one long timestep) in Figure 1.
Figure 1: Illustration of stuttering transitions (in yellow) and spurious trajectories (in red and blue). In yellow: the timestep is too small, and therefore the reachable set intersects with the starting cell. In red: from the initial green cell, two cells are reachable with a jump of length $\tau$. At the next step, the two reachable sets from both cells are computed, and intersect with 12 cells. Applying the same input for $2\tau$ is discarded because 3 of the 12 cells intersect with an obstacle. In blue: contrarily to the red case, we directly consider a timestep corresponding to a larger $2\tau$ timestep. The reachable set intersects with 4 cells that do not intersect with the obstacle. It is therefore valid to apply the same input for $2\tau$ from the initial green cell.
This conservativism leads either to infeasible problems, or to sub-optimal solutions in the context of optimal control. On the other hand, considering small timesteps allows for more flexibility, and is sometimes needed in order to ensure feasibility of the problem.
@@
In this work, we propose to take into account the tradeoff between optimality and feasibility by introducing temporal abstractions, in which both small and large timesteps are allowed. We prove the correctness of such abstractions by proving that the latter are in $\operatorname{FRR}$ with the concrete continuous-time systems, thereby providing formal guarantees. We then provide the necessary theoretical framework to use these abstractions in the context of optimal control under reachability specifications. Finally, we demonstrate that our method outperforms abstraction techniques based on fixed time-step discretizations by yielding better controllers.
@@
The sets $\mathbb{R},\mathbb{Z},\mathbb{Z}_{+}$ denote respectively the sets of real numbers, integers and non-negative integer numbers. For example, we use $[a,b]\subseteq\mathbb{R}$ to denote a closed continuous interval and $[a;b]=[a,b]\cap\mathbb{Z}$ for discrete intervals. The symbol $\varnothing$ denotes the empty set. Given two sets $\mathcal{X}_{1},\mathcal{X}_{2}$, we define a single-valued map as $f:\mathcal{X}_{1}\rightarrow\mathcal{X}_{2}$, while a set-valued map is defined as $F:\mathcal{X}_{1}\rightarrow 2^{\mathcal{X}_{2}}$, where $2^{\mathcal{X}_{2}}$ is the power set of $\mathcal{X}_{2}$, i.e., the set of all subsets of $\mathcal{X}_{2}$. The image of a subset $\Omega\subseteq\mathcal{X}_{1}$ under $F:\mathcal{X}_{1}\rightarrow 2^{\mathcal{X}_{2}}$ is denoted $F(\Omega)$. Given two sets $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, we identify a binary relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ with set-valued maps, i.e., $R(x_{1})=\{x_{2}\mid(x_{1},x_{2})\in R\}$ and $R^{-1}(x_{2})=\{x_{1}\mid(x_{1},x_{2})\in R\}$. A relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ is said to be strict if $R(x_{1})\neq\varnothing$ for all $x_{1}\in\mathcal{X}_{1}$. When $\mathcal{X}_{2}$ is finite, the mapping $R:\mathcal{X}_{1}\to 2^{\mathcal{X}_{2}}$ is referred to as a quantizer. The set of maps $\mathcal{A}\rightarrow\mathcal{B}$ is denoted $\mathcal{B}^{\mathcal{A}}$, and the set of all signals that take their values in $\mathcal{B}$ and are defined on intervals of the form $[0;N)=[0;N-1]$ is denoted $\mathcal{B}^{\infty}$, $\mathcal{B}^{\infty}=\bigcup_{N\in\mathbb{N}\cup\{\infty\}}\mathcal{B}^{[0;N)}$. Given a set-valued map $f:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ and $\boldsymbol{x}\in\mathcal{X}^{[0;N)}$, we denote by $f(\boldsymbol{x})=\{\boldsymbol{y}\in\mathcal{Y}^{[0;N)}\mid\forall k\in[0;N):\ y(k)\in f(x(k))\}$.
@@ II-A Control framework
The use of a set-valued function to describe the transition map of a system allows to model perturbations and any other kind of non-determinism in a unified formalism. We say that a transition system is deterministic if for every state $x\in\mathcal{X}$ and input $u\in\mathcal{U}$, $F(x,u)$ is either empty or a singleton. Otherwise, we say that it is non-deterministic. A finite transition system, in contrast to an infinite system, refers to a system characterized by finitely many states and inputs.
We define the behavior of a system as its set of maximal trajectories.
@@
$\mathcal{B}(\mathcal{S})=\{(\boldsymbol{x},\boldsymbol{u})\mid(\boldsymbol{x},\boldsymbol{u})\text{ is a trajectory of }\mathcal{S}\text{ on }[0;N),\\ \text{ and if }N<\infty\text{ then }F(x(N-1),u(N-1))=\varnothing\},$ (2)
is called the behavior of $\mathcal{S}$. $\triangle$
@@
This controller is static because the allowed inputs at each state depend only on the current state, not on the past trajectory.
We now define the control problem.
@@ II-B Abstraction-based control
Proposition 1 ([18, Theorem V.4])
Let $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ and $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$ be transition systems, and let $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$. If $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$, then for any controller $\mathcal{C}_{2}$, the controller $\mathcal{C}_{1}=\mathcal{C}_{2}\circ R$ ensures
$\mathcal{B}(\mathcal{C}_{1}\times\mathcal{S}_{1})\subseteq R^{-1}(\mathcal{B}(\mathcal{C}_{2}\times\mathcal{S}_{2})).$
Consequently, if $\Sigma_{1}\subseteq R^{-1}(\Sigma_{2})$ and $\mathcal{C}_{2}$ solves $(\mathcal{S}_{2},\Sigma_{2})$, then $\mathcal{C}_{1}$ solves $(\mathcal{S}_{1},\Sigma_{1})$.
Although finiteness of $\mathcal{S}_{2}$ is not required, in practice, leveraging powerful graph-theoretical tools requires the construction of a finite abstraction of the underlying dynamical system for the abstract controller synthesis (see examples in Section IV).
@@ III-A Temporal abstractions
First, any continuous-time system $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$ can be framed as a transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ as follows:
@@
Now, in this work, we apply abstraction-based control to the continuous-time system $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$ (see Definition 6). Consequently, this approach involves constructing both a transition system $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$ and a relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ such that $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$, where $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ is the transition system associated with $\mathcal{D}$ as defined in (6). As proven below, the following conditions on the abstract system $\mathcal{S}_{2}$ and $R$ are sufficient for $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$ to hold:
@@
Let $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$ be a continuous-time system with its associated transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ as in (6). Consider a transition system $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$ and a strict relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$, with $\mathcal{U}_{2}$ as in (7), and $R$ and $F_{2}$ as in (8). Then it holds that $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$.
@@ III-B Application to optimal control
We begin by defining the considered reach-avoid specification (see Definition 4). Note that our method can be generalized to arbitrary specifications $\Sigma$, including those defined using linear temporal logic (LTL) [4, Chapter 5] but, for the sake of clarity, we leave this extension for further work.
@@
In addition to specifying the state behavior of the closed-loop system, transition costs can be incorporated to address optimal control problems aiming to minimize the worst-case cost.
@@
$[B_{c}(v)](x)=\min_{u\in\mathcal{U}_{\mathcal{S}}(x)}\left(c(x,u)+\max_{x^{+}\in F(x,u)}v(x^{+})\right).$ (12)
@@
is commonly referred to as a Bellman value function.
While the Bellman value function $v$, which satisfies the Bellman equation $v(x)=[B_{c}(v)](x)$ (13) allows to implicitly characterize the optimal solution, its computation is generally intractable for infinite state systems. Instead, we consider suboptimal value functions [16, Definition 7], which satisfy the relaxed inequality $v(x)\geq[B_{c}(v)](x)$.
@@
Then, the controller $\mathcal{C}$ associated with $v$ solves the control problem $(\mathcal{S},[\operatorname{dom}(\mathcal{C}),\mathcal{F},\mathcal{O}])$. If $\mathcal{I}\subseteq\operatorname{dom}(\mathcal{C})$, then $\mathcal{C}$ also solves $(\mathcal{S},\Sigma^{\text{Reach}})$. Moreover, the value function $v(x)$ provides an upper bound on the worst-case cost under $\mathcal{C}$, i.e., $v(x)\geq\ell_{\mathcal{C}}(x)$, and the overall performance satisfies $\mathcal{L}(\mathcal{C})\geq\mathcal{L}(\mathcal{C}^{*})$, with equality if $v$ is a Bellman value function.
@@
In the special case where $R$ is deterministic, the value function defined in (15) reduces to $v_{1}=v_{2}\circ R$, as previously established in [16, Theorem 8].
In practice, while the finiteness of $\mathcal{S}_{2}$ enables efficient computation of a Bellman value function $v_{2}$ using classical graph-theoretic tools such as dynamic programming [6, Chapter 1], Theorem 2 allows it to be translated into a suboptimal value function $v_{1}$ for the concrete system.

### v2

@@ Abstract
As control systems grow in complexity, abstraction-based methods have become essential for designing controllers with formal guarantees. However, a key limitation of these methods is their reliance on discrete-time models, typically obtained by discretizing continuous-time systems with a fixed timestep. This discretization leads to two major problems: when the timestep is small, the abstraction includes numerous stuttering and spurious trajectories, making controller synthesis suboptimal or even infeasible; conversely, a large time step may also render control design infeasible due to a lack of flexibility. In this work, drawing inspiration from Reinforcement Learning concepts, we introduce temporal abstractions, which allow for a flexible timestep. We provide a method for constructing such abstractions and formally establish their correctness in controller design. Furthermore we show how to apply these to optimal control under reachability specifications. Finally we showcase our methods on two numerical examples, highlighting that our approach leads to controllers that achieve a lower worst-case control cost.
@@ I INTRODUCTION
The growing complexity of engineered dynamical systems has made controller design increasingly arduous. In the literature, this paradigm shift has been coined by the academic and industrial communities as the cyber-physical revolution [12, 3, 14]. On the other hand, it is crucial to provide safety and/or performance guarantees for such systems [13]. In particular, we focus on reachability analysis [2]: designing a controller that drives the system from an initial set to a target set while avoiding obstacles. Being able to solve such problems for complex dynamical systems is crucial in many applications (see [1, 8] for applications to autonomous driving and robotic aerial vehicles).
In this context, abstraction-based control, also known as symbolic control, is a promising tool to provide formal reachability guarantees [19]. Every abstraction-based control method follows the same systematic procedure:
@@
    Abstraction – The concrete system is approximated by a finite-state mathematical object (the abstract system);
@@
    Concretization – From the discrete controller, a concrete controller is refined for the original (concrete) dynamical system.
@@
Continuous-time deterministic dynamical systems are abstracted by partitioning the state space into a finite set of cells, selecting a finite set of inputs, considering discrete-time dynamics with timestep $\tau$. The abstract states correspond to the cells, and a transition between $s_{1}$ and $s_{2}$ with input $u$ is added if the set of reachable states from the cell $s_{1}$ intersects with the cell $s_{2}$. In this work, we focus on the time discretization. Indeed, in the context of reachability analysis, fixing a too small constant timestep $\tau$ leads to conservatism for at least two reasons:
@@
    Spurious trajectories – At each step of a trajectory in the abstract system, the information of where exactly the state is within the cell is lost. As a consequence, an error is propagated with the number of steps, and more and more trajectories are valid in the abstraction but cannot happen in reality: such trajectories are called spurious trajectories. This phenomenon is also illustrated in red (two small timesteps) and in blue (one long timestep) in Figure 1.
Figure 1: Illustration of stuttering transitions (yellow) and spurious trajectories (red and blue). Yellow: the timestep is too small, and therefore the reachable set intersects with the starting cell. Red: from the initial green cell, two cells are reachable with a jump of length $\tau$. At the next step, the two reachable sets from both cells are computed. Applying the same input for $2\tau$ is discarded because 3 cells intersect with the black obstacle. Blue: we now directly consider a timestep corresponding to a larger timestep, $2\tau$. The reachable set intersects with 4 cells that do not intersect with the obstacle, making the input valid.
This conservatism leads either to infeasible problems, or to sub-optimal solutions in the context of optimal control. On the other hand, considering small timesteps allows for more flexibility, and is sometimes needed in order to ensure feasibility of the problem.
@@
In this work, we propose to take into account this trade-off by introducing temporal abstractions, in which multiple timesteps are allowed. We prove the correctness of such abstractions, thereby providing formal guarantees. We then provide the necessary theoretical framework to use these abstractions in the context of optimal control under reachability specifications. While our approach does not directly tackle the curse of dimensionality inherent to abstraction-based control, numerical experiments show that using multiple timesteps yields lower worst-case cost controllers than those obtained with a finer input space discretization.
@@
Given two sets $\mathcal{X}_{1},\mathcal{X}_{2}$, we define a single-valued map as $f:\mathcal{X}_{1}\rightarrow\mathcal{X}_{2}$, while a set-valued map is defined as $F:\mathcal{X}_{1}\rightarrow 2^{\mathcal{X}_{2}}$, where $2^{\mathcal{X}_{2}}$ is the power set of $\mathcal{X}_{2}$, i.e., the set of all subsets of $\mathcal{X}_{2}$. The image of a subset $\Omega\subseteq\mathcal{X}_{1}$ under $F:\mathcal{X}_{1}\rightarrow 2^{\mathcal{X}_{2}}$ is denoted $F(\Omega)$. Given two sets $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, we identify a binary relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ with set-valued maps, i.e., $R(x_{1})=\{x_{2}\mid(x_{1},x_{2})\in R\}$ and $R^{-1}(x_{2})=\{x_{1}\mid(x_{1},x_{2})\in R\}$. A relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ is said to be strict if $R(x_{1})\neq\varnothing$ for all $x_{1}\in\mathcal{X}_{1}$. When $\mathcal{X}_{2}$ is finite, the mapping $R:\mathcal{X}_{1}\to 2^{\mathcal{X}_{2}}$ is referred to as a quantizer. The set of all signals that take their values in $\mathcal{B}$ are defined on intervals of the form [0;N)=[0;N1][0;N)=[0;N-1] is denoted $\mathcal{B}^{\infty}$, $\mathcal{B}^{\infty}=\bigcup_{N\in\mathbb{N}\cup\{\infty\}}\mathcal{B}^{[0;N)}$. Given a set-valued map $f:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ and $\boldsymbol{x}\in\mathcal{X}^{[0;N)}$, we denote by $f(\boldsymbol{x})=\{\boldsymbol{y}\in\mathcal{Y}^{[0;N)}\mid\forall k\in[0;N):\ y(k)\in f(x(k))\}$.
@@ II-A Control framework
We use a set-valued function to model perturbations and non-determinism in a unified formalism. We say that the transition system is deterministic if for every state $x\in\mathcal{X}$ and input $u\in\mathcal{U}$, $F(x,u)$ is either empty or a singleton
A finite transition system refers to a system characterized by finitely many states and inputs. We define the behaviour of a system as its set of maximal trajectories.
@@
[…]
is called the behaviour of $\mathcal{S}$. $\triangle$
@@
This controller is static because the allowed inputs at each state depend only on the current state, not on the past trajectory. We now define the control problem.
@@ II-B Abstraction-based control
Proposition 1 ([17, Theorem V.4])
Let $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ and $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$ be transition systems, and let $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$. If $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$, then for any controller $\mathcal{C}_{2}$, the controller $\mathcal{C}_{1}=\mathcal{C}_{2}\circ R$ ensures $\mathcal{B}(\mathcal{C}_{1}\times\mathcal{S}_{1})\subseteq R^{-1}(\mathcal{B}(\mathcal{C}_{2}\times\mathcal{S}_{2}))$. Consequently, if $\Sigma_{1}\subseteq R^{-1}(\Sigma_{2})$ and $\mathcal{C}_{2}$ solves $(\mathcal{S}_{2},\Sigma_{2})$, then $\mathcal{C}_{1}$ solves $(\mathcal{S}_{1},\Sigma_{1})$.
Altough $\mathcal{S}_{2}$ need not be finite, practical use of graph-theoretic tools for controller synthesis requires a finite abstraction (see Section IV).
@@ III-A Temporal abstractions
We begin by introducing a general framework to frame a continuous-time system as a transition system without loss of generality, by extending the input space:
Definition 7
Consider a continuous-time system $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$. This system can be framed as a transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ as follows:
@@
We call this system the multi-sampled system. $\triangle$
Now, in this work, we apply abstraction-based control to the continuous-time system $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$ (see Definition 6). Consequently, this approach involves constructing both a transition system $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$ and a relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$ such that $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$, where $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ is the transition system associated with $\mathcal{D}$ as in Definition 7. As proven below, the following conditions on the abstract system $\mathcal{S}_{2}$ and $R$ are sufficient for $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$ to hold:
@@
Let $\mathcal{D}=(\mathcal{X},\mathcal{U},f)$ be a continuous-time system with its associated transition system $\mathcal{S}_{1}=(\mathcal{X}_{1},\mathcal{U}_{1},F_{1})$ as in (6). Consider a transition system $\mathcal{S}_{2}=(\mathcal{X}_{2},\mathcal{U}_{2},F_{2})$ and a strict relation $R\subseteq\mathcal{X}_{1}\times\mathcal{X}_{2}$, with $\mathcal{U}_{2}$ as in (7), and $R$ and $F_{2}$ as in (8). Then it holds that $\mathcal{S}_{1}\preceq_{R}^{\operatorname{FRR}}\mathcal{S}_{2}$. $\triangle$
@@ III-B Application to optimal control
We begin by defining the considered reach-avoid specification (see Definition 4).
@@
In addition to specifying the state behaviour of the closed-loop system, transition costs can be incorporated to address optimal control problems aiming to minimize the worst-case cost.
@@
$B_{c}(v)](x)=\min_{u\in\mathcal{U}_{\mathcal{S}}(x)}\left(c(x,u)+\max_{x^{+}\in F(x,u)}v(x^{+})\right).$ (12)
@@
is commonly referred to as a Bellman value function. Its computation is generally intractable for infinite state systems. Instead, we consider suboptimal value functions [15, Definition 7].
@@
Then, the controller $\mathcal{C}$ associated with $v$ solves the control problem $(\mathcal{S},[\operatorname{dom}(\mathcal{C}),\mathcal{F},\mathcal{O}])$. If $\mathcal{I}\subseteq\operatorname{dom}(\mathcal{C})$, then $\mathcal{C}$ also solves $(\mathcal{S},\Sigma^{\text{Reach}})$. Moreover, the value function $v(x)$ provides an upper bound on the worst-case cost under $\mathcal{C}$, i.e., $v(x)\geq\ell_{\mathcal{C}}(x)$, and the overall performance satisfies $\mathcal{L}(\mathcal{C})\geq\mathcal{L}(\mathcal{C}^{*})$, with equality if $v$ is a Bellman value function. $\triangle$
@@
In the special case where $R$ is deterministic, the value function defined in (15) reduces to $v_{1}=v_{2}\circ R$, as previously established in [15, Theorem 8]. This is the case in the classical abstraction approach [17, 19], where the abstract state space corresponds to a partition of (a subset of) the concrete state space. In practice, while the finiteness of $\mathcal{S}_{2}$ enables efficient computation of a Bellman value function $v_{2}$ using classical graph-theoretic tools such as dynamic programming [5, Chapter 1], Theorem 2 allows it to be translated into a suboptimal value function $v_{1}$ for the concrete system.